In an experiment designed to measure the heat of fusion of ice, 25 grams of ice at 0

oC is


dropped into 200 grams of water at 30


oC contained in a copper calorimeter of mass 100 g. The


final temperature is 17.5


oC. What value of the heat of fusion do this value yield? Specific


heat of copper = 0.0931 cal/gC


o

. (10


1
Expert's answer
2022-02-19T06:55:01-0500

Let, mass of water used =m Since, Heat gained = Heat lostm1​cΔt1​=mL+m2​cΔt2​m×4.2×(60−10)=(40×336)+(40×4.2×(10−0))m×4.2×50=(40×336)+1680Therefore, the mass of water used, m=42×5
13440+1680​=210
15120​=72g